solve by elimination method: 9x-9y=9;
7y-3x=-5

Answer by ewatrrr(24785)


Hi
solving by elimination method:
9x-9y= 9
7y-3x=-5

9x- 9y =9
-9x +21y = -15
12y = -6
y = -(1/2)
solving for x
7y-3x=-5
-(7/2) + 5 = 3x
3/2 = 3x
1/2 = x
CHECKING our answer
9x-9y= 9/2 - 9(-1/2) = 9/2 + 9/2 = 9